(​Risk-adjusted

NPV​)

The Hokie Corporation is considering two mutually exclusive projects. Both require an initial outlay of

​$13,000

and will operate for

9

years. Project A will produce expected cash flows of

​$4,000

per year for years 1 through

9

whereas project B will produce expected cash flows of

​$5,000

per year for years 1 through

9.

Because project B is the riskier of the two​ projects, the management of Hokie Corporation has decided to apply a required rate of return of

18

percent to its evaluation but only a required rate of return

11

percent to project A. Determine each​ project's risk-adjusted net present value.

What is the​ risk-adjusted NPV of project​ A?
Answer #1

Project A

NPV= -Initial cost + PV of future cash flows

PV of future cash flows = PV of annuity = Annuity*(1-1/(1+rate)^number of terms)/rate

= 4000*(1-1/1.11^9)/0.11

=22148.19

NPV = -13000+22148.19

= $9148.19

Project B

NPV= -Initial cost + PV of future cash flows

PV of future cash flows = PV of annuity = Annuity*(1-1/(1+rate)^number of terms)/rate

= 5000*(1-1/1.18^9)/0.18

=21515.11

NPV = -13000+21515.11

=$8515.11